OpenStudy (anonymous):

Solve the equation. 6x – 9 = 5x – 4

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

subtract 5x from both sides and add 9 to both sides

myininaya (myininaya):

6x-5x=-4+9 x=5

OpenStudy (anonymous):

6x – 9 = 5x – 4 6x-5x=9-4 x=5